Equal weights of methane and hydrogen are mixed in an empty container at 25°C. The fraction of total pressure exerted by hydrogen is
1. 1/2 .                2.8/9 .           3. 16/19.       4.1/9

Grade:9

1 Answers

Arun
5 years ago
Let x grams of each hydrogen and methane are mixed, Moles of H2 = x/2 Moles of CH4 = x/16 ⇒ Mole-fraction of H2 = (x/2)/(x/2 + x/16) = 8/9 ⇒ (Partial pressure of H_2)/(Total pressure) = mole-fraction of H2 = 8/9